Let y^2=4ax be a parabola and p be a point on it. Let the normal at p intersects the x-axis at Q. Draw a line at Q perpendicular to the above normal. Show that the line intersects the parabola y^2+4a(x-2a)=0
1
Expert's answer
2020-06-22T18:19:54-0400

let us consider the point P(a,2a) on the parabola "y^2=4ax."

The rough diagram of the question is attached here...




from P the normal to the parabola "y^2=4ax" is drawn which intersects the x axis at the point Q,the coordinate of which is (3a,0). ( working of Q is shown below).

from Q a line perpendicular to PQ is drawn and we have to show that it intersects the parabola

"y^2+4a(x-2a)=0."


Now equation of tangent at the point P(a,2a) is given by

"y.2a=2a(x+a)\\or \space y=x+a \space \space ..............(1)" (Here slope (m) of tangent is 1.)

"\therefore" slope of the Normal is "-1."

Now Equation of Normal (which is perpendicular to the tangent at P ) is given by

"(y-2a)=-1(x-a)\\or \space y=-x+3a.\\"

AS this normal intersects X-axis at Q. so ordinate of Q must be '0'.

putting y=0 in "y=-x+3a."

we get x=3a.

"\therefore" The coordinate of Q is (3a,0).


Now slope of the line which is passing through Q (3a,0) and perpendicular to the normal at P

is again "1" as product of slope of two perpendiculars line is always "-1" .


"\therefore" The equation of line passing through Q(3a,0) is given by

"(y-0)=1(x-3a)\\or \space y=x-3a."

Now the coordinates of the end points of the latus rectum of the parabola "y^2=4ax \space is \\"

"P(a,2a) \space and (a,-2a)."


clearly the Coordinate (a,-2a) satisfies both "y=x-3a \space and \space y^2+4a(x-2a)=0."

Hence the required line "(y=x-3a)" intersects the parabola "y^2+4a(x-2a)=0."

(PROVED).